Tom invests $18,750 in two seperate accounts. One earns 10% simple interest and the other 12%. If he earned a total of $2117, how much did he invest at 10%?
Thanks!

Answer by rfer(16322)

.12x+.10(18750-x)=2117
.12x+1875-.10x=2117
.02x=242
x=$12,100 at 12%
18750-12100=$6,650 at 10%